Find all solutions of the equation.
tan^2 = 1/3
Thanks so much, after someone hopefully answers this, I will do the rest of my homework around this example :)

Answer by ewatrrr(24785)

tanx = sqrt(1/3) = 1/ √3
(cos, sinx) summary for Unit Circle